'n' drops of same size are charged to potential V volts each. if ther are added up to form a single drop then its potential will be.?

Grade:

1 Answers

Ashmita Bhattacharyya
11 years ago

For 1 drop, V=kq/r..(r=radius of each drop)

therefore,q=Vr/k....for 1 drop

if we have n drops, total charge=Q=nq=Vrn/k

total radius=R=rn^.33...(apply volume conservation)

V=kQ/R=kVrn/k(rn^.33)=Vn^.66(Ans.)

Hope this helps.